OpenStudy (anonymous):

Let u = [-1,1,2] and v= [5,-3,2]. Find the equation of the plane through origin that consist of all linear combinations of u and v.

OpenStudy (anonymous):

Start with the given equation. Plug in the given vectors and then Multiply the scalars by EVERY element in the vectors after multiplying Add the vectors by adding the corresponding components.

OpenStudy (anonymous):

What is the given equation? ax + by+ cz = 0?

OpenStudy (anonymous):

First take any two linear combinations of \(\vec{u}\) and \(\vec{v}\). I used \(\vec{u}+\vec{v}\) and \(\vec{u}-\vec{v}\) and the answer seems to work. \[\vec{u}+\vec{v}=\langle4,-2,4\rangle\\\vec{u}-\vec{v}=\langle-6,4,0\rangle\] Find the normal vector, which is given by the cross product of any two vectors in the plane. These are two such vectors, so the normal vector is \[\vec{n}=\begin{vmatrix}\vec{i}&\vec{j}&\vec{k}\\4&-2&4\\-6&4&0\end{vmatrix}=\langle a,b,c\rangle\] \(a,b,c\) correspond to the coefficients in the plane equation, \(ax+by+cz=0\).

OpenStudy (anonymous):

so would a, b, and c be equal to whatever I get for i, j, and k?

OpenStudy (amistre64):

1u + 0v and 0u + 1v would also have worked :)

OpenStudy (amistre64):

the ijk are what we 'solve' the psuedo determinant for. in other words, the cross product gives us:\[a\hat i+b\hat j+c\hat k\]which represents the vector (a,b,c)

OpenStudy (anonymous):

I worked it out and solved for the vector (a,b,c) to be (-4,12,-2). Do I then put it into an equation or is -4x + 12y -2z = d my final answer?

OpenStudy (amistre64):

not quite. we want to anchor this vector to one of the given points: (x1,y1,z1) such that a(x-x1)+b(y-y1)+c(z-z1)=0 this will then expand to some: ax + by + cz = d

OpenStudy (amistre64):

-(5z-6x-2y) x y z x y [-1,1,2] -1 1 [5,-3,2] 5 -3 +(2x+10y+3z) +6x +2y -5z ------------ 8 12 -2 4,6,-1 is what i get, but lets chk with the wolf

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%5B-1%2C1%2C2%5D+cross+%5B5%2C-3%2C2%5D 8,12,-2 so a scaled version of it is 4,6,-1 is fine since its the direction that counts and not the length

OpenStudy (amistre64):

siths cross gets us: (-16, -24, 4), which reduces to 4,6,-1 as well when we scale it by -1/4

OpenStudy (anonymous):

I don't understand where you got the -(5z-6x-2y) from.

OpenStudy (anonymous):

is that from x-x1?

OpenStudy (amistre64):

i did a quick method of a 3x3 determinant is all

OpenStudy (amistre64):

show me how you crossed it .... since its the crossing result thats important and not the method used to cross.

OpenStudy (anonymous):

How I crossed u and v?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

if you wanna learn the shortcut, its simple enough to do and easy to remember if you know the 2x2 shortcut

OpenStudy (anonymous):

I did i j k -1 1 2 5 -3 2 and crossed them to get -4i + 12j - 2k

OpenStudy (amistre64):

yeah, your end result is not correct .... which means you prolly had a simple error in the process

OpenStudy (amistre64):

spose to get 8i not -4i

OpenStudy (anonymous):

I just realized that as you posted. So does that give me a normal? I don't understand at all what you're telling me.

OpenStudy (amistre64):

i j k -1 1 2 5 -3 2 i * 1 2 = i|2-(-6)| = 8i -3 2

OpenStudy (amistre64):

the crossing gives you a normal vector to both u and v, yes

OpenStudy (amistre64):

lets use the point (-1,1,2) as an anchor for the normal: 4(x+1)+6(y-1)-(z-2)=0 this is fine and dandy for a plane equation, but they tend to like to see it in its expanded form soo 4x +6y -z + (6-6+2) = 0 4x +6y -z +2 = 0 or 4x +6y -z = -2

OpenStudy (amistre64):

hmm, typo 4(1) is not 6 lol

OpenStudy (amistre64):

pfft, been too long. the constant will be zero since its thru the origin to start with

OpenStudy (anonymous):

So what do you do with the -2? I understand how you got the equation now, just I don't understand what you're supposed to do with the -2

OpenStudy (anonymous):

does it just disappear?

OpenStudy (amistre64):

see it for the typo that it is and correct it :)

OpenStudy (anonymous):

Oh okay! I see it now

OpenStudy (amistre64):

4(x+1)+6(y-1)-(z-2)=0 4x+4+6y-6-z+2=0 :) 4-6+2 = 0 is your constant, not -2

OpenStudy (amistre64):

now if we hadnt reduced the magnitude of the normal we would have gotten 8x + 12y -2z = 0 which is not in 'simplest' form since its just a multiple of what we have.

OpenStudy (anonymous):

but either answer will work correct?

OpenStudy (amistre64):

depends on whos grading it and what they expect from you.

OpenStudy (amistre64):

any positive, nonzero multiple of (4x+6y-z=0) will be the equation of the plane, but its up to your own material how specific it needs to get

OpenStudy (anonymous):

I don't imagine they'll mark it wrong if it's in its simplest form. I have this question solved but I have a few others that I might just need a nudge in the right direction for, if I need help I will post it

OpenStudy (amistre64):

very well :)